An element crystallizes in a body centered cubic ubit cell. The raduis of this element is 2.2x10^-10 m, and the density is 0.991 g/cm^3. How many moles of atoms are within one unit cell?
1
Expert's answer
2020-12-16T05:52:28-0500

"6.02\u00d710" 23"Units\equiv" 1 mole

Body centered cubic has 2 units

"([\frac{1}{8}×8]+1)"

Using proportions,

2 units will be "\frac{2units×1mole}{\left(6.02×10^{\smash{23}}\right) units}=\left(3.22×10^{\smash{-24}}\right)moles"